a class has 24 boys and 16 girls. On a test the class mean was 75. the mean of the girl's score was 72. What was the mean of the boy's sccores?
Answer by mananth(12270)

a class has 24 boys and 16 girls. On a test the class mean was 75. the mean of the girl's score was 72. What was the mean of the boy's sccores?
24+16 = 40 students.... mean = 75
Total scored by 40 students = 3000
Girls = 16 ....their mean = 72
total scored by girls = 16*72= 1152
Difference between the scores will give total score of boys.
3000-1152 = 1848
Number of boys = 24
their mean will be 1848 / 24
=77